Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP:; phone: +1 575-835-7370


News, Travel, Visitors (McMullin)

  • Overall status: In A configuration (08 Jun); Correlator hardware (64 BlB) all available. Sundry antenna-based issues (some limitations from spare parts). Push on CBE stability and improved data rates.
    • Banff SKA meeting (Myers): 2-day science meeting; 1 day public forum (quite political); 1 day industry meeting - pointed questions about SKA. Site decision on the SKA soon (Moran is chair of this).
  • Key dates:
    • 08 Jun - 12 Sep: A configuration (move from 31 May - 08 Jun).
    • 01 Aug - Proposal Deadline (C, CnB, B)
  • Staff Travel:
    • RP: 10 Jul - 23 Jul (UK, SA)
    • RP: 09 Aug - 21 Aug (URSI)
    • FO: 12 Aug - 12 Sep
    • RP: 10 Sep - 16 Sep (Brazil)
  • Visitors:
    • Robert Harris (AC982): 24 Apr - 16 Jul (High frequency w/ Claire)
    • Michiel Brentjens: 11 Jun - 11 Nov (Holography; characterizing polarized beams)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA21 back to array today? EA23 next
    • Receivers: Ku for EA21 this week.
    • Calibration (Perley): Gain compression/expansion still being looked at. (at the site today). Antennas with new RF switches; tests run on this morning.
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): Stable weekend; problems last night still under investigation. Push on data rates
      • Fast-dump status (Rupen/Law): 27 antennas (64 ch) ~ 300 ms dumps = 6 MB/s; looked at the crab and did 7 antennas with 12 ms dumps (10 MB/s sustainable. Still exploring further.
  • Computing (Robnett):
    • Notes: New 2TB disks have started arriving.

Software status and updates (Butler)

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Configuration move week (work around these activities:
  • Mon
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Tue
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday): Use for targeted testing/science
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues.Identified possible issues in the noise related to the 3-level phase rotator.
  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags). .Needs close-out testing.
  • Correlator: [Clark/Carlson] Delay rocking across bands. Ongoing discussion of implementation; pending Bryan's return
  • Array: [Gustaaf] Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).
  • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week/soon

  • CM/CBE: Data rate testing and push; work on new missing scans.
  • Holography mode testing.
  • V Stokes follow-up.
  • Subarrays: cleanup.
  • 3-bit sampler testing; automatic gain setting.
  • Stage 2 mixer testing to relax sub 128 MHz BW grids
  • HPC large dataset (500 GB)
  • Executor fshift algorithm (new; increments <100 Hz)
  • Phased array/astrometry mode testing.
  • Correlator: Rupen: S-band shadow testing
  • Spectral line: Momjian: HI High resolution bandpass testing
  • Baseband edge/filter rolloff.
  • Sub-band dependent fshifts.
  • Array: *; Individual issue sleuthing (focus on phase jumps).

Commissioning Groups

    • 4-band (Owen):
      • Notes:
    • Adv: Low band (Owen):
      • Notes:
    • Adv: Astrometry (Mioduszewski):
      • Notes: Begin testing of UVW precision issues (Brunthaler). Report next week.
  • Calibrator Models
    • Notes: A configuration data being collected.
  • HPC (Bhatnagar/Owen/Golap):
    • Notes:


RSRO A configuration status:

Individual reports (updates)

  • Irwin:


  • RSRO Calendar:
  • RSRO DnC, C config status:
  • Test Project Codes == *TRSR*
  • Guidelines for RSRO observing:
  • Notes:
    • Current Residents
      • D.J. Saikia (10C-119/AI145)::RSRO-75; 300G(2); x7106-POC: Momjian (Low Frequency)
      • J. Irwin (10C-119AI145)::RSRO-75; 300G(1); x7333 -- POC: Myers (Low Frequency)
      • Peter Hofner (10B-124/AH1011)::RSRO-26; 200G(1); x7496 -- POC: Ott (CASA testing)
      • Jacqueline van Gorkom (10C-211/AV324); 200D/x7098 -- POC: Momjian (Low Frequency); Not present
      • Andreas Brunthaler::200E/x7246 -- POC: Mioduszewski (Astrometry/VLBI)
      • Sergio Dzib::200G/x7462 -- POC: Mioduszewski (Astrometry)
      • Jane Greaves::(AC982)::RSRO-4; 300G;x7106 -- POC: Chandler (High Frequency Documentation for Workshop)
      • Satoko Takahashi:: 200D/x7098--POC: Ott (Imaging)

Demo Science Status

Attached is an image of w50 made with new EVLA C array data combined
with D array data (observed by Gloria et al)

I have mosaiced the 2 arrays independently and used the mean psf of
the D and C mosaics to 'feather' the images together.

I have not yet done a full D+C array joint deconvolution; the problem
being the weights between VLA and EVLA data.

Natural weighting was used just to get maximum Sig/Noise.

Still some artifact of calibration / interference can be seen as low
level ripples or stripes in places

Some improvements obviously still to come...once i figure the weight
transfer and some of the low level interference plguing some fields.


  • Exposure calculator and Java - anyone else with issues; found wrong Java version was being used by default.

Misc Actions

  • Old/Low priority
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • [Rupen]: Development of WIDAR guidelines document

-- JosephMcMullin - 2011-07-11
  • w50_cd.image.jpg:
    w50 cd.image.jpg
Topic revision: r3 - 2011-07-19, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback